# Inferring CSM Properties of Type II SNe Using a Magnitude-Limited ZTF Sample

### v2

@@ 2.5.2 Rise Time Limitations
Figure 5: Phase-space diagram showing peak absolute magnitude in ZTF $r$ ($M_{r,peak}$) vs. rest-frame duration above half-maximum brightness ($t_{50}$) for all SNe with $m_{peak}\leq 19$ mag classified in the BTS through December 31, 2023. Type II SNe are colour-coded by subtype, with other SN classes (e.g. Ia, Ib/c, SLSNe-I) shown in grey for comparison. $M_{r,peak}$ corrected for Galactic extinction only.
@@ 3.1 Overall Distribution
Studies of the relationship between rise time and peak luminosity in Type II SNe have yielded conflicting results. Significant correlations have been reported for Type II, IIb, and IIn SNe (e.g. González-Gaitán et al., 2015; Pessi et al., 2019; Hiramatsu et al., 2024), suggesting the rise and peak may be intrinsically coupled by their power source. However, other analyses find no significant correlation (e.g. Gall et al., 2015; Rubin et al., 2016; Valenti et al., 2016; Nyholm et al., 2020). These discrepant findings likely stem from small sample numbers, limiting their statistical power to comment of population characteristics.
Motivated by the possibility of an enhanced population with fast rises and luminous peaks, either from SBO occurring in the CSM or at the surface of the star, we tested the strength of any existing correlation between $t_{25\_75}$ and $M_{g,peak}$. For Type II SNe, a Spearman rank test between $t_{25\_75}$ in ZTF $g$ and $M_{g,peak}$ finds a correlation, with a correlation coefficient $\rho=-0.21$ and p-value $p<10^{-5}$. The scatter present in the correlation is likely caused by the large diversity of Type II SNe and lack of clear division within Type II SNe (e.g. IIP vs IIL vs 87A-like) as from Fig. 4, it can be seen that Type II SNe occupy both bright and faint, fast and slow regions, see Fig. 3.
In our sample of 110 Type IIn SNe (which includes 16 SNe classified as super-luminous), we find $\rho=-0.18$ and p-value $p\sim 0.05$, a correlation both weaker and less significant than the recent findings of Hiramatsu et al. (2024). We see great diversity in our Type IIn light curve morphology, which seems to suggest a range of progenitor pathways are possible, with a large range in $t_{25\_75}$ of $\sim$ 2 d to 40 d and $M_{g,peak}$ of $-22.20$ to $-17.01$ mag. As Type IIn SNe are well-understood to be CSM-driven (e.g. Schlegel, 1990; Fassia et al., 2000; Smartt, 2009; Smith et al., 2010; Taddia et al., 2013; Ransome et al., 2021), and under this scenario it is expected that both the rise time and luminosity increase with the amount of CSM present (continuing the interaction), this possible correlation is not surprising (Section section 6). For Type IIb SNe, we find no significant correlation as $p>0.1$.
@@ 3.2 Volume Corrected Distributions
To represent the intrinsic distribution of parameters, we use Kernel Density Estimation (KDE). For each observed value, we generate a normalised Gaussian kernel centred on that real value and weight each Gaussian as $1/V_{max}$. The width (sigma) of each kernel is optimised using cross-validation (e.g. Wu, 1997). This approach estimates the underlying probability density function of the parameter distribution. We also normalise the KDE by the sum of weights, allowing us to account for Malmquist bias.
@@ 4 Methods II – Comparison to Simulated Light Curves
Light curve models are increasingly incorporating CSM or extended stellar envelopes (e.g. Hillier & Dessart, 2012; Dessart et al., 2015, 2017; Das & Ray, 2017; Morozova et al., 2018; Pearson et al., 2023; Tinyanont et al., 2022; Moriya et al., 2023; Morag et al., 2023). Morozova et al. (2018) demonstrated that including CSM in SNEC  [^9] Supernova ExplosioN Code. (Morozova et al., 2015) models significantly improved fits for 20 well-observed, multi-band light curves to estimate progenitor parameters, such as $R_{CSM}$, CSM density, and $M_{CSM}$ (see Figs. 2 and 3 and Table 2 in Morozova et al., 2018). Similar conclusions emerge from studies by Das & Ray (2017), Bruch et al. (2023), Moriya et al. (2023), Irani et al. (2024), and Jacobson-Galán et al. (2024a), which collectively find that substantial CSM masses near Type II SN progenitors are common and deposited shortly before core-collapse.
@@ 4.1 Measuring Theoretical Light Curve Metrics
We measure various rise times – including the same $t_{25\_75}$ described in Section section 2.5 – directly from the M23 ZTF bandpass light curves, in addition to absolute magnitudes (at peak and various $N$ days after peak), magnitude decline rates and colours, e.g. $(g-r)_{g,max}$ – Fig. 1 shows example measurements using ZTF SN light curves.
@@
In Fig. 8, we see a clear bi-modality, suggesting the dichotomy seen in observations is reflecting the transition between purely shock cooling (SC) dominated rises and rises dominated by the interaction heating from SBO shocking the CSM (e.g. Irani et al., 2024; Jacobson-Galán et al., 2024a). Correlation tests between progenitor-SN parameters confirm this dichotomy stems from significant relationships between CSM properties and light curve observables, as well as between different CSM parameters [^12] Correlations between CSM parameters likely reflect physics pre-defined in the simulations.. This is further evidenced by the colour gradient seen when we apply a colour map based on $\dot{M}$  or $R_{CSM}$ – see section A.5 for further details.
Two distinct populations emerge in the theoretical models: fast risers ($t_{25\_75}$leq 5 d) with moderately more luminous peaks, possessing higher $M_{CSM}$ and smaller $R_{CSM}$ (e.g. confined and dense); and slower risers ($t_{25\_75}$ $>$ 5 d) with overall less luminous peaks, less massive $M_{CSM}$ and larger $R_{CSM}$ (e.g. less confined and less dense). Within the slower population, the most luminous events still require substantial $M_{CSM}$, suggesting CSM mass remains a key driver of peak luminosity. This bi-modality emerges naturally from the underlying physics rather than parameter choices, hinting at fundamental differences in mass-loss mechanisms.
For slower-rising events, the correlation between $t_{25\_75}$ and CSM parameters lessens, replaced by a stronger dependence on progenitor mass ($M_{ZAMS}$; which serves as a proxy for progenitor radius). This shift likely reflects both the increasing dominance of light travel time and diffusion processes in more massive, extended progenitors, as well as the slower cooling of the shocked envelope (e.g. Nakar & Sari, 2010; Ofek et al., 2010; Tominaga et al., 2011; Morozova et al., 2016; Irani et al., 2024).
@@ 5 Physical Progenitor Property Inference Analysis
To quantify the percentage of RSGs that are surrounded by significant components of CSM at the time of core-collapse, we use the early light curve to infer the presence and properties of CSM. As the rise of Type II SNe is highly sensitive to the CSM parameters and progenitor radius (e.g. Morozova et al., 2016; Morozova et al., 2018; Pearson et al., 2023; Tinyanont et al., 2022; Moriya et al., 2023; Irani et al., 2024), we can estimate and place constraints on these properties from photometry alone. The high cadence and good coverage on the rise specifically, combined with the high completeness of the BTS lends itself well to such a detailed study.
@@ 5.1 Defining Light Curve - CSM Relations
The performance of each polynomial order is analysed using the Bayesian Information Criterion (BIC; Table 5). We split the correlation testing into $t_{25\_75}$leq 5 d and $>$ 5 d as the correlation appears strongest for the fast-rising models $t_{25\_75}$leq 5 d and under-predicts $M_{CSM}$ for several long-rising $t_{25\_75}$ $>$ 5 d models by up to 1 – 2 orders of magnitude – see Fig. 10. This likely reflects the physics of CSM interaction: confined, dense CSM shells enable a rapid and efficient conversion of kinetic energy to radiation, producing brief, bright emission with fast rise times (e.g. Moriya et al., 2011; Chevalier, 2012; Tinyanont et al., 2022; Pearson et al., 2023; Li et al., 2024). In contrast, longer rise times can result from either more extended CSM configurations where energy is released more gradually, or from different physical processes entirely, weakening the direct correlation between $M_{CSM}$ and light curve properties (e.g. Moriya, 2023; Jacobson-Galán et al., 2024a).
@@
$R_{CSM}$ measurements become most reliable in regimes where $M_{CSM}$ is large enough to influence observables such as $M_{g,peak}$ and $t_{25\_75}$ significantly (e.g. make $M_{g,peak}$ brighter or $t_{25\_75}$ shorter). We determine the critical value – below which the CSM does not notably influence the early light curve – of $M_{CSM}$ by varying the physical parameters $\dot{M}$, $R_{CSM}$, and $\beta$ independently (noting that $M_{CSM}$ is a function of these three variables within the model). Below this minimum $M_{CSM}$, the CSM will become too diffuse to meaningfully influence the early light curve evolution, effectively transitioning to a regime where CSM interaction is negligible. This physical expectation is reinforced by the distinct bi-modality observed in M23’s theoretical models,
@@ 5.2 Mass-Loss Rate
For Type II SNe, we assume a stellar wind velocity of 10 km s-1, consistent with previous literature (e.g. Davies et al., 2022; Moriya et al., 2023). We also calculate the time in which the mass was removed $t_{removal}\sim R_{CSM}$/$v_{wind}$. While M23 explored $R_{CSM}$ from $10^{14}$ – $10^{15}$ cm, we cannot uniquely determine $\dot{M}$  from our observations alone. For comparison with previous studies, we adopt a fiducial value of $R_{CSM}$ = 5×$\times 10^{14}$ cm.
@@ 5.3 Progenitor Property Volume Corrected Distributions
For comparison, recent studies of SN 2023ixf (e.g. Hosseinzadeh et al., 2023; Hiramatsu et al., 2023; Jencson et al., 2023; Jacobson-Galán et al., 2023; Li et al., 2024; Zimmerman et al., 2024) and SN 2024ggi (e.g. Jacobson-Galán et al., 2024b; Shrestha et al., 2024; Chen et al., 2024a, b) prefer $\dot{M}$  between $10^{-3}$ – $10^{-2}$ $\textup{M}_{\odot}$ yr-1 calculated by photometric and/or spectroscopic modelling of the event and its environment. For SN 2023ixf, we find $M_{CSM}$ $\sim$ 0.1 $\textup{M}_{\odot}$, $R_{CSM}$ $\sim$ 6×$\times 10^{14}$ cm and $\dot{M}$ $\sim$ 1×$\times 10^{-2}$ $\textup{M}_{\odot}$ yr-1 respectively. Singh et al. (2024) and Moriya & Singh (2024) conduct a similar exercise, finding the best fitting M23 models to SN 2023ixf, and measure similar values for $\dot{M}$, $10^{-3}$ –$10^{-2}$ $\textup{M}_{\odot}$ yr-1 and $R_{CSM}$ $\sim$ 5–10×$\times 10^{14}$ cm. For SN 2024ggi, we find $M_{CSM}$ $\sim$ 5×$\times 10^{-3}$ $\textup{M}_{\odot}$, $R_{CSM}$ $\sim$ 3×$\times 10^{14}$ cm and $\dot{M}$ $\sim$ 2×$\times 10^{-4}$ $\textup{M}_{\odot}$ yr-1 respectively.
@@ 6.1 Luminosity and Rise Distributions
We have limited overlap with Pessi et al. (2024a), studying Type II SLSNe from ZTF, though we find $23^{+9}_{-7}$ype IIn SNe in our weighted sample have $M_{g,peak}$ brighter than $-19$ mag – this is the most appropriate as we grouped Type IIn and SLSNe together – with a large range in $M_{g,peak}$ $-17.01$ to $-22.20$ mag. Our unweighted median $M_{g,peak}$ for Type IIn, $-19.32~{}\pm~{}0.13$ mag, agrees with median values in Hiramatsu et al. (2024).
@@
Type II SNe exhibit diverse rise times, with their distribution (Fig. 4) suggesting a continuous rather than bi-modal range of progenitor properties. The lack of bi-modality in the observed $t_{25\_75}$ distributions is significant and likely reflects an overabundance of these events compared to their representation in the M23 grid. This continuous distribution is seen before and after the volume correction, Fig. 7(a), suggesting this is not a result of lacking observations of a particular population. The prevalence of events with both rapid rise times and high luminosities provides compelling evidence for dense CSM being common among Type II SNe. These characteristics are consistent with SBO within CSM (e.g. Moriya et al., 2011; Chevalier, 2012; Das & Ray, 2017; Moriya et al., 2018; Tinyanont et al., 2022; Pearson et al., 2023; Li et al., 2024), as models without significant CSM struggle to simultaneously produce such fast rise times and enhanced peak brightness.

### v3

@@ 2.5.2 Rise Time Limitations
Figure 5: Phase-space diagram showing peak absolute magnitude in ZTF $r$ ($M_{r,peak}$) vs. rest-frame duration above half-maximum brightness ($t_{50}$) for all SNe with $m_{peak}\leq 19$ mag classified in the BTS through December 31, 2023. Type II SNe are colour-coded by subtype, with other SN classes (e.g., Ia, Ib/c, SLSNe-I) shown in grey for comparison. $M_{r,peak}$ corrected for Galactic extinction only.
@@ 3.1 Overall Distribution
Studies of the relationship between rise time and peak luminosity in Type II SNe have yielded conflicting results. Significant correlations have been reported for Type II, IIb, and IIn SNe (e.g., González-Gaitán et al., 2015; Pessi et al., 2019; Hiramatsu et al., 2024), suggesting the rise and peak may be intrinsically coupled by their power source. However, other analyses find no significant correlation (e.g., Gall et al., 2015; Rubin et al., 2016; Valenti et al., 2016; Nyholm et al., 2020). These discrepant findings likely stem from small sample numbers, limiting their statistical power to comment of population characteristics.
Motivated by the possibility of an enhanced population with fast rises and luminous peaks, either from SBO occurring in the CSM or at the surface of the star, we tested the strength of any existing correlation between $t_{25\_75}$ and $M_{g,peak}$. For Type II SNe, a Spearman rank test between $t_{25\_75}$ in ZTF $g$ and $M_{g,peak}$ finds a correlation, with a correlation coefficient $\rho=-0.21$ and p-value $p<10^{-5}$. The scatter present in the correlation is likely caused by the large diversity of Type II SNe and lack of clear division within Type II SNe (e.g., IIP vs IIL vs 87A-like) as from Fig. 4, it can be seen that Type II SNe occupy both bright and faint, fast and slow regions, see Fig. 3.
In our sample of 110 Type IIn SNe (which includes 16 SNe classified as super-luminous), we find $\rho=-0.18$ and p-value $p\sim 0.05$, a correlation both weaker and less significant than the recent findings of Hiramatsu et al. (2024). We see great diversity in our Type IIn light curve morphology, which seems to suggest a range of progenitor pathways are possible, with a large range in $t_{25\_75}$ of $\sim$ 2 d to 40 d and $M_{g,peak}$ of $-22.20$ to $-17.01$ mag. As Type IIn SNe are well-understood to be CSM-driven (e.g., Schlegel, 1990; Fassia et al., 2000; Smartt, 2009; Smith et al., 2010; Taddia et al., 2013; Ransome et al., 2021), and under this scenario it is expected that both the rise time and luminosity increase with the amount of CSM present (continuing the interaction), this possible correlation is not surprising (Section section 6). For Type IIb SNe, we find no significant correlation as $p>0.1$.
@@ 3.2 Volume Corrected Distributions
To represent the intrinsic distribution of parameters, we use Kernel Density Estimation (KDE). For each observed value, we generate a normalised Gaussian kernel centred on that real value and weight each Gaussian as $1/V_{max}$. The width (sigma) of each kernel is optimised using cross-validation (e.g., Wu, 1997). This approach estimates the underlying probability density function of the parameter distribution. We also normalise the KDE by the sum of weights, allowing us to account for Malmquist bias.
@@ 4 Methods II – Comparison to Simulated Light Curves
Light curve models are increasingly incorporating CSM or extended stellar envelopes (e.g., Hillier & Dessart, 2012; Dessart et al., 2015, 2017; Das & Ray, 2017; Morozova et al., 2018; Pearson et al., 2023; Tinyanont et al., 2022; Moriya et al., 2023; Morag et al., 2023). Morozova et al. (2018) demonstrated that including CSM in SNEC  [^9] Supernova ExplosioN Code. (Morozova et al., 2015) models significantly improved fits for 20 well-observed, multi-band light curves to estimate progenitor parameters, such as $R_{CSM}$, CSM density, and $M_{CSM}$ (see Figs. 2 and 3 and Table 2 in Morozova et al., 2018). Similar conclusions emerge from studies by Das & Ray (2017), Bruch et al. (2023), Moriya et al. (2023), Irani et al. (2024), and Jacobson-Galán et al. (2024a), which collectively find that substantial CSM masses near Type II SN progenitors are common and deposited shortly before core-collapse.
@@ 4.1 Measuring Theoretical Light Curve Metrics
We measure various rise times – including the same $t_{25\_75}$ described in Section section 2.5 – directly from the M23 ZTF bandpass light curves, in addition to absolute magnitudes (at peak and various $N$ days after peak), magnitude decline rates and colours, e.g., $(g-r)_{g,max}$ – Fig. 1 shows example measurements using ZTF SN light curves.
@@
In Fig. 8, we see a clear bi-modality, suggesting the dichotomy seen in observations is reflecting the transition between purely shock cooling (SC) dominated rises and rises dominated by the interaction heating from SBO shocking the CSM (e.g., Irani et al., 2024; Jacobson-Galán et al., 2024a). Correlation tests between progenitor-SN parameters confirm this dichotomy stems from significant relationships between CSM properties and light curve observables, as well as between different CSM parameters [^12] Correlations between CSM parameters likely reflect physics pre-defined in the simulations.. This is further evidenced by the colour gradient seen when we apply a colour map based on $\dot{M}$  or $R_{CSM}$ – see section A.5 for further details.
Two distinct populations emerge in the theoretical models: fast risers ($t_{25\_75}$leq 5 d) with moderately more luminous peaks, possessing higher $M_{CSM}$ and smaller $R_{CSM}$ (e.g., confined and dense); and slower risers ($t_{25\_75}$ $>$ 5 d) with overall less luminous peaks, less massive $M_{CSM}$ and larger $R_{CSM}$ (e.g., less confined and less dense). Within the slower population, the most luminous events still require substantial $M_{CSM}$, suggesting CSM mass remains a key driver of peak luminosity. This bi-modality emerges naturally from the underlying physics rather than parameter choices, hinting at fundamental differences in mass-loss mechanisms.
For slower-rising events, the correlation between $t_{25\_75}$ and CSM parameters lessens, giving way to a stronger dependence on progenitor mass ($M_{ZAMS}$), which serves as a proxy for progenitor radius. This transition reflects both the increasing dominance of light travel time and diffusion processes in more massive, extended progenitors, as well as the slower cooling of the shocked envelope (e.g., Nakar & Sari, 2010; Ofek et al., 2010; Tominaga et al., 2011; Morozova et al., 2016; Irani et al., 2024). Notably, for events with longer rise times ($\geq$ 5 d), the progenitor radius influences rise time by affecting the SBO time and location (Chevalier & Irwin, 2011; Moriya et al., 2011; Svirski et al., 2012; González-Gaitán et al., 2015; Morozova et al., 2016), though the light curve evolution is still predominantly shaped by CSM interaction rather than envelope cooling, depending critically on the CSM density profile (Moriya et al., 2023; Irani et al., 2024).
@@ 5 Physical Progenitor Property Inference Analysis
To quantify the percentage of RSGs that are surrounded by significant components of CSM at the time of core-collapse, we use the early light curve to infer the presence and properties of CSM. As the rise of Type II SNe is highly sensitive to the CSM parameters and progenitor radius (e.g., Morozova et al., 2016; Morozova et al., 2018; Pearson et al., 2023; Tinyanont et al., 2022; Moriya et al., 2023; Irani et al., 2024), we can estimate and place constraints on these properties from photometry alone. The high cadence and good coverage on the rise specifically, combined with the high completeness of the BTS lends itself well to such a detailed study.
@@ 5.1 Defining Light Curve - CSM Relations
The performance of each polynomial order is analysed using the Bayesian Information Criterion (BIC; Table 5). We split the correlation testing into $t_{25\_75}$leq 5 d and $>$ 5 d as the correlation appears strongest for the fast-rising models $t_{25\_75}$leq 5 d and under-predicts $M_{CSM}$ for several long-rising $t_{25\_75}$ $>$ 5 d models by up to 1 – 2 orders of magnitude – see Fig. 10. This likely reflects the physics of CSM interaction: confined, dense CSM shells enable a rapid and efficient conversion of kinetic energy to radiation, producing brief, bright emission with fast rise times (e.g., Moriya et al., 2011; Chevalier, 2012; Tinyanont et al., 2022; Pearson et al., 2023; Li et al., 2024). In contrast, longer rise times can result from either more extended CSM configurations where energy is released more gradually, or from different physical processes entirely, weakening the direct correlation between $M_{CSM}$ and light curve properties (e.g., Moriya, 2023; Jacobson-Galán et al., 2024a).
@@
$R_{CSM}$ measurements become most reliable in regimes where $M_{CSM}$ is large enough to influence observables such as $M_{g,peak}$ and $t_{25\_75}$ significantly (e.g., make $M_{g,peak}$ brighter or $t_{25\_75}$ shorter). We determine the critical value – below which the CSM does not notably influence the early light curve – of $M_{CSM}$ by varying the physical parameters $\dot{M}$, $R_{CSM}$, and $\beta$ independently (noting that $M_{CSM}$ is a function of these three variables within the model). Below this minimum $M_{CSM}$, the CSM will become too diffuse to meaningfully influence the early light curve evolution, effectively transitioning to a regime where CSM interaction is negligible. This physical expectation is reinforced by the distinct bi-modality observed in M23’s theoretical models,
@@ 5.2 Mass-Loss Rate
For Type II SNe, we assume a stellar wind velocity of 10 km s-1, consistent with previous literature (e.g., Davies et al., 2022; Moriya et al., 2023). We also calculate the time in which the mass was removed $t_{removal}\sim R_{CSM}$/$v_{wind}$. While M23 explored $R_{CSM}$ from $10^{14}$ – $10^{15}$ cm, we cannot uniquely determine $\dot{M}$  from our observations alone. For comparison with previous studies, we adopt a fiducial value of $R_{CSM}$ = 5×$\times 10^{14}$ cm.
@@ 5.3 Progenitor Property Volume Corrected Distributions
For comparison, recent studies of SN 2023ixf (e.g., Hosseinzadeh et al., 2023; Bostroem et al., 2023; Hiramatsu et al., 2023; Jencson et al., 2023; Jacobson-Galán et al., 2023; Li et al., 2024; Zimmerman et al., 2024) and SN 2024ggi (e.g., Jacobson-Galán et al., 2024b; Shrestha et al., 2024; Chen et al., 2025, 2024) prefer $\dot{M}$  between $10^{-3}$ – $10^{-2}$ $\textup{M}_{\odot}$ yr-1 calculated by photometric and/or spectroscopic modelling of the event and its environment. For SN 2023ixf, we find $M_{CSM}$ $\sim$ 0.1 $\textup{M}_{\odot}$, $R_{CSM}$ $\sim$ 6×$\times 10^{14}$ cm and $\dot{M}$ $\sim$ 1×$\times 10^{-2}$ $\textup{M}_{\odot}$ yr-1 respectively. Singh et al. (2024) and Moriya & Singh (2024) conduct a similar exercise, finding the best fitting M23 models to SN 2023ixf, and measure similar values for $\dot{M}$, $10^{-3}$ –$10^{-2}$ $\textup{M}_{\odot}$ yr-1 and $R_{CSM}$ $\sim$ 5–10×$\times 10^{14}$ cm. For SN 2024ggi, we find $M_{CSM}$ $\sim$ 5×$\times 10^{-3}$ $\textup{M}_{\odot}$, $R_{CSM}$ $\sim$ 3×$\times 10^{14}$ cm and $\dot{M}$ $\sim$ 2×$\times 10^{-4}$ $\textup{M}_{\odot}$ yr-1 respectively.
@@ 6.1 Luminosity and Rise Distributions
We have limited overlap with Pessi et al. (2025), studying Type II SLSNe from ZTF, though we find $23^{+9}_{-7}$ype IIn SNe in our weighted sample have $M_{g,peak}$ brighter than $-19$ mag – this is the most appropriate as we grouped Type IIn and SLSNe together – with a large range in $M_{g,peak}$ $-17.01$ to $-22.20$ mag. Our unweighted median $M_{g,peak}$ for Type IIn, $-19.32~{}\pm~{}0.13$ mag, agrees with median values in Hiramatsu et al. (2024).
@@
Type II SNe exhibit diverse rise times, with their distribution (Fig. 4) suggesting a continuous rather than bi-modal range of progenitor properties. The lack of bi-modality in the observed $t_{25\_75}$ distributions is significant and likely reflects an overabundance of these events compared to their representation in the M23 grid. This continuous distribution is seen before and after the volume correction, Fig. 7(a), suggesting this is not a result of lacking observations of a particular population. The prevalence of events with both rapid rise times and high luminosities provides compelling evidence for dense CSM being common among Type II SNe. These characteristics are consistent with SBO within CSM (e.g., Moriya et al., 2011; Chevalier, 2012; Das & Ray, 2017; Moriya et al., 2018; Tinyanont et al., 2022; Pearson et al., 2023; Li et al., 2024), as models without significant CSM struggle to simultaneously produce such fast rise times and enhanced peak brightness.